Please help...
One number is 8 less than another number. The sum of four times the larger and twice the smaller is 20. Find the larger number.
I got the second equation as 4n+8(2)=20 with n=1 but I don't know how to use that to find the answer.

Answer by RAY100(1637)

Let n = one number,,,,,,,other number = (n-8)
.
4(n) +2(n-8) = 20
.
4n +2n -16 = 20
.
6n = 36
.
n=6,,,,,,(n-8) = -2,,,,,,numbers
.
.
check,,,4 (6) + 2(-2) = 20,,,,,ok
.